OpenStudy (anonymous):

A rectangular garden is 18 feet wide. Its length is 8 feet more than twice the width. What is the area of the garden?

OpenStudy (lucaz):

area of rectangle is width * length, the length is the width +8, so you have \[\Large{w\times l=18\times (18+8)=468~ft^2}\]

OpenStudy (anonymous):

8 feet more than twice the width....

OpenStudy (lucaz):

oh, sorry, than the (18+8) should be (2*18)+8

OpenStudy (lucaz):

\[\Large{w\times l=18\times (2\times 18)+8)=792~ft^2}\]

OpenStudy (anonymous):

Yes Thanks!!

OpenStudy (lucaz):

no problem